# Continuous Approximations of Projected Dynamical Systems via Control Barrier Functions

### v1

@@ IV The relationship between PDSs and CBF-based dynamics
In this section, we delve into the relationship between PDS-based dynamics (8) and CBF-based dynamics (10). First, we prove that CBF-based dynamics are perturbations of PDSs. Then, building on this, we prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDSs as $\alpha\to\infty$. The proofs to our results can be found in Section VII, unless otherwise stated.
@@ IV-A CBF-based dynamics as perturbations of PDSs
$M_{1}:=\min_{x\in\partial\mathcal{S}}\|\nabla h(x)\|,\quad M_{2}:=\max_{x\in\partial\mathcal{S}}\|\nabla h(x)\|,\quad M_{3}:=M_{2}+L_{\nabla h}\gamma(\frac{1}{\alpha_{*}}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|),$
@@
Prop. IV.1 extends [19, Thm. III.1] in two ways. First, towards addressing interconnections (8) and (10), it considers $f:{}^{m}\times{}^{n}\to{}^{n}$, in contrast to [19, Thm. III.1], which only considers $f:{}^{n}\to{}^{n}$. Second, it gives bounds $\delta$ on the normal cone elements in the right-hand side. This is crucial for proving convergence of trajectories later on, as it enables results from [25] for well-posed DIs, requiring local boundedness of the set-valued map. normal-$\blacksquare$
Observe that, due to Assumption 1 item 3, $M_{1}>0$, and, hence, there is always an $\epsilon$ such that $0<\epsilon<\min_{x\in\partial\mathcal{S}}\|\nabla h(x)\|$. Moreover, both $\sigma(\alpha)$ and $\delta$ are well-defined, as they are maxima of continuous locally bounded functions over compact sets. Prop. IV.1 gives rise to the two following results:
@@
$\begin{bmatrix}\dot{\zeta}\\ \dot{\xi}\end{bmatrix}\in\begin{bmatrix}g(\zeta,\xi)\\ K_{\sigma(\alpha)}\bigg{(}\mathrm{N}^{P}_{\mathcal{S}}\Big{(}\xi,f(\zeta,\xi),\delta\Big{)}\bigg{)}\end{bmatrix},(\zeta,\xi)\in{}^{m}\times(\mathcal{S}+\sigma(\alpha)\mathbb{B})$ (11)
@@
Let us explain how Cors. IV.1, IV.2 and IV.3 are integral to our work. First, note that solutions of the CBF-based dynamics (10) are solutions of perturbed DI (11) (from Cor. IV.1). Also, notice that $\sigma$ is continuous, strictly decreasing on $\alpha$, and satisfies $\sigma(\alpha)\geq 0$ and $\lim_{\alpha\to\infty}\sigma(\alpha)=0$. Thus, from Cor. IV.2, as $\alpha\to\infty$, the perturbed DI (11), and thus the CBF-based dynamics (10), tends to the limiting DI (12); and, the larger $\alpha$ is picked, the closer is perturbed DI (11), and thus the CBF-based dynamics (10), to the limiting DI (12). Finally, from Cor. IV.3, we know that the limiting DI (12) shares the same solutions with the PDS (7). Thus, CBF-based dynamics (10) are perturbed versions of PDS-based ones (7), and $\sigma(\alpha)$ is a bound on that perturbation, that vanishes as $\alpha\to\infty$. In the coming section, we employ all the above, to prove that trajectories of CBF-based dynamics (10) uniformly converge to trajectories of PDS-based dynamics (7).
@@
[20, Prop. 4.4] proves that the right-hand sides of (7) and (10) coincide at $\alpha\to\infty$. Nonetheless, this is generally not enough to infer convergence of solutions. In contrast, here we provide quantitative bounds, depending on $\alpha$, on the perturbation that (10) is to (7). As becomes evident from the following section, this is what is indeed needed to prove convergence of solutions. normal-$\blacksquare$
Remark 4.
Notice that we only study solutions that are constrained in the compact set $\mathcal{Z}\times\mathcal{S}$. On the other hand, $\mathcal{Z}\times\mathcal{S}$ can be chosen arbitrarily large. In this sense, our results are semiglobal. normal-$\blacksquare$
@@ IV-B Trajectories of CBF-based dynamics converge to trajectories of PDSs
Here, we employ Cors. IV.1, IV.2 and IV.3, to show that trajectories of CBF-based dynamics (10) uniformly converge to bounded trajectories of PDSs (7), (8), as $\alpha\to\infty$. This establishes that CBF-based dynamics are legitimate approximations of PDSs.
Specifically, we first show that any sequence of bounded solutions to the CBF-based dynamics, corresponding to an increasing sequence $\{\alpha_{i}\}_{i\in\mathbb{N}}$ of parameters $\alpha_{i}$ with $\lim_{i\to\infty}\alpha_{i}=\infty$, has a subsequence which converges to a solution of the PDS (7). Moreover, the whole sequence of bounded solutions to the CBF-based dynamics converges to a solution of the PDS (7), when this solution is unique.
@@
Once Props. IV.1 and VII.1 (See Section VII), and Cors. IV.1, IV.2 and IV.3 are given, the proof of Thm. IV.1 follows steps similar to [10, Section 4]. normal-$\blacksquare$
While Thm. IV.1 establishes uniform convergence of solutions of CBF-based dynamics (10) to solutions of PDSs (7), it assumes that solutions of CBF-based dynamics stay in $\mathcal{Z}\times\mathcal{S}$. And, while it is guaranteed that the $\xi$-part of the solution is always in $\mathcal{S}$, this is not always the case for the $\zeta$-part. Nevertheless, if PDS (7) obeys a stability property, then, as the following result shows, we can guarantee that solutions to CBF-based dynamics stay in $\mathcal{Z}\times\mathcal{S}$, for sufficiently large $\alpha$, and thus they converge to solutions of PDS (7).
@@
Consider vector fields $f:{}^{m}\times{}^{n}\to{}^{n}$ and $g:{}^{m}\times{}^{n}\to{}^{m}$, a compact set $\mathcal{Z}\subset{}^{m}$ and a set $\mathcal{S}\subset{}^{n}$ defined by (1). Let Assumptions 1 and 2 hold. Further, let $g$ be continuous and locally bounded. Given $\gamma>0$, consider compact sets $\mathcal{Z}_{0},\mathcal{Z}^{\prime}$, such that $\mathcal{Z}_{0}\subseteq\mathcal{Z}^{\prime}$ and $\mathcal{Z}^{\prime}+\gamma\mathbb{B}\subseteq\mathcal{Z}$. Given any initial condition $(z_{0},x_{0})\in\mathcal{Z}_{0}\times\mathcal{S}$ and a $T>0$, assume any solution $\phi:[0,T]\to{}^{m}\times{}^{n}$ to PDS (7), with $\phi(0)=(z_{0},x_{0})$, satisfies $\phi(t)\in\mathcal{Z}^{\prime}\times\mathcal{S}$ for all $t\in[0,T]$. Then, there exists $\alpha^{\prime}\geq\alpha_{*}$, such that for any $\alpha\geq\alpha^{\prime}$, the solution $\psi:[0,T]\to{}^{m}\times{}^{n}$ to the CBF-based dynamics (10), with $\psi(0)=(z_{0},x_{0})$, satisfies $\psi(t)\in\mathcal{Z}\times\mathcal{S}$ for all $t\in[0,T]$. Therefore, for any increasing sequence $\{\alpha_{i}\}_{i\in\mathbb{N}}$, such that $\alpha_{i}\geq\alpha^{\prime}$ and $\lim_{i\to\infty}\alpha_{i}=\infty$, and respective sequence of solutions $\{\psi_{i}\}_{i\in\mathbb{N}}$ to CBF-based dynamics (10), statements 1 and 2 of Thm. IV.1 hold.
@@
The stability assumption of Thm. IV.2 is not strong, for practical considerations. When PDSs are introduced as feedback controllers [8, 10, 5, 9, 7, 11], this is done in a way that guarantees stability of the closed-loop implementation (7). Thus, if the PDS-based controller is replaced by a CBF-based one, to yield the CBF-based dynamics (10), Thm. IV.2 ensures that trajectories of (10) converge to trajectories of (7), for $\alpha\geq\alpha^{\prime}$; that is, (10) behaves similarly to (7). normal-$\blacksquare$
Remark 7 (Stability of CBF-based dynamics).
Combining Theorem IV.2 and [25, Prop. 6.34], one can derive a stability result for CBF-based dynamics: if the origin for PDS-based dynamics (7) is asymptotically stable, then it is practically stable for CBF-based dynamics (10). Asymptotic stability and contractivity results for CBF-based dynamics have also been reported in [21]. normal-$\blacksquare$
V Numerical Example: Feedback Optimization
@@
$A=\begin{bmatrix}-1&1\\ 0&-2\end{bmatrix},\text{ }B=\begin{bmatrix}0\\ 1\end{bmatrix},\text{ }\Phi(\zeta)=\Big{(}\zeta-\begin{bmatrix}1\\ 1\end{bmatrix}\Big{)}^{2},\text{ }\mathcal{S}=\{u\in\real:0.36-u^{2}\geq 0\}$
@@
(a) Evolution of the control input
(b) Evolution of state $\zeta_{1}$
(c) Evolution of state $\zeta_{2}$
@@ VI Conclusion
We have proven that trajectories of CBF-based dynamics uniformly converge to trajectories of PDS-based ones, as the tunable parameter $\alpha$ of the CBF-constraint is taken to $\infty$. Towards this, we have also proven that CBF-based dynamics are perturbations of PDSs, and quantified the perturbation. One can employ our results to implement discontinuous PDS-based controllers in a continuous manner, employing CBFs. We have demonstrated this on a numerical example emanating from feedback optimization. Similarly, our results enable novel numerical simulation schemes for PDSs, through CBF-based dynamics, overcoming disadvantages of existing schemes, such as projections to possibly non-convex sets. Finally, the established bridge between PDSs and CBFs might yield other results in the future, such as insights on stability.
VII Technical results and proofs
VII-A Proof of Prop. IV.1
To prove Prop. IV.1, we make use of Lemmas VII.1, VII.2 and VII.3 below. These lemmas have already appeared in [19], in a simpler form; here they are extended, to account for the fact that we study interconnections (8) and (10), rather than autonomous setups (6) and (9), which was the context of [19]. [^5] Here, $f:{}^{m}\times{}^{n}\to{}^{n}$, that is $f$ has not only $\xi$ as an argument, but $\zeta$ too; see (6) and (9). Moreover, in [19] the proofs of Lemmas VII.1 and VII.2 were only sketched. Here, we provide the full proofs of Lemmas VII.1 and VII.2 and recall Lemma VII.3.
The following results are stated under the assumptions of Prop. IV.1. Given any $z\in\mathcal{Z}$, we denote $U_{\mathrm{cbf},\alpha}(z):=\{x\in\mathcal{S}\mid\mathcal{L}_{f}h(z,x)+\alpha h(x)\leq 0\}$.
Lemma VII.1.
Given $\alpha>0$ and $z\in\mathcal{Z}$, consider any $x\in U_{\mathrm{cbf},\alpha}(z)$ and any $y\in\mathrm{proj}_{\partial\mathcal{S}}(x)$. It holds that
$\|x-y\|\leq\gamma(\frac{1}{\alpha}|\mathcal{L}_{f}h(z,x)|)$
Proof of Lemma VII.1.

### v3

@@ IV The relationship between PDSs and CBF-based dynamics
In this section, we delve into the relationship between PDS-based dynamics (8) and CBF-based dynamics (10). First, we prove that CBF-based dynamics are perturbations of PDSs. Then, building on this, we prove that trajectories of CBF-based dynamics uniformly converge to trajectories of PDSs as $\alpha\to\infty$. The proofs to our results can be found in Section VIII, unless otherwise stated.
@@ IV-A CBF-based dynamics as perturbations of PDSs
$M_{1}:=\min_{x\in\partial\mathcal{S}}\|\nabla h(x)\|,\quad M_{2}:=\max_{x\in\partial\mathcal{S}}\|\nabla h(x)\|,\quad M_{3}:=M_{2}+L_{\nabla h}\gamma(\frac{1}{\alpha_{*}}\max_{(z,x)\in\mathcal{Z}\times\mathcal{S}}|\mathcal{L}_{f}h(z,x)|)$
@@
Prop. IV.1 extends [19, Thm. III.1] in two ways. First, towards addressing interconnections (8) and (10), it considers $f:^{m}\times^{n}\to^{n}$, in contrast to [19, Thm. III.1], which only considers $f:^{n}\to^{n}$. Second, it gives bounds $\delta$ on the normal cone elements in the right-hand side. This is crucial for proving convergence of trajectories later on, as it enables results from [25] for well-posed DIs, requiring local boundedness of the set-valued map. $\blacksquare$
Observe that, due to Assumption 1 item 3, $M_{1}>0$, and, hence, there is always an $\epsilon$ such that $0<\epsilon<\min_{x\in\partial\mathcal{S}}\|\nabla h(x)\|$. Moreover, both $\sigma(\alpha)$ and $\delta$ are well-defined, as they are maxima of continuous locally bounded functions over compact sets. The set $\mathcal{Z}$ constrains the first argument of $f$, i.e., $z\in\mathcal{Z}\subset^{m}$. We thus study solutions of PDS-based dynamics (7) and CBF-based dynamics (10) that are constrained in $\mathcal{Z}\times\mathcal{S}$, as it becomes clear below (see Remark 6).
Prop. IV.1 gives rise to the two following results:
@@
$\begin{bmatrix}\dot{\zeta}\\ \dot{\xi}\end{bmatrix}\in\begin{bmatrix}g(\zeta,\xi)\\ K_{\sigma(\alpha)}\bigg{(}\mathrm{N}^{P}_{\mathcal{S}}\Big{(}\xi,f(\zeta,\xi),\delta\Big{)}\bigg{)}\end{bmatrix},\quad(\zeta,\xi)\in^{m}\times(\mathcal{S}+\sigma(\alpha)\mathbb{B})$ (11)
@@
Let us explain the significance of Cors. IV.1, IV.2 and IV.3. First, note that solutions of the CBF-based dynamics (10) are solutions of perturbed DI (11) (from Cor. IV.1). Also, notice that $\sigma$ is continuous, strictly decreasing on $\alpha$, and satisfies $\sigma(\alpha)\geq 0$ and $\lim_{\alpha\to\infty}\sigma(\alpha)=0$. Thus, from Cor. IV.2, as $\alpha\to\infty$, the perturbed DI (11), and thus the CBF-based dynamics (10), tends to the limiting DI (12); and, the larger $\alpha$ is picked, the closer is perturbed DI (11), and thus the CBF-based dynamics (10), to the limiting DI (12). Finally, from Cor. IV.3, we know that the limiting DI (12) shares the same solutions with the PDS (7). Thus, CBF-based dynamics (10) are perturbed versions of PDS-based ones (7), and $\sigma(\alpha)$ is a bound on that perturbation, that vanishes as $\alpha\to\infty$. In the coming section, we employ all the above, to prove that trajectories of CBF-based dynamics (10) uniformly converge to trajectories of PDS-based dynamics (7).
@@
[20, Prop. 4.4] proves that the right-hand sides of (7) and (10) coincide at $\alpha\to\infty$. This is generally not enough to infer convergence of solutions. In contrast, we provide bounds, depending on $\alpha$, on the perturbation that (10) is to (7). As it becomes evident in Section IV-B, this enables proving convergence of solutions. $\blacksquare$
@@ IV-B Trajectories of CBF-based dynamics converge to trajectories of PDSs
Here, we employ Cors. IV.1, IV.2 and IV.3, to show that trajectories of CBF-based dynamics (10) uniformly converge to bounded trajectories of PDSs (7), (8), as $\alpha\to\infty$. This establishes that CBF-based dynamics are indeed approximations of PDSs. Specifically, we first show that any sequence of bounded solutions to the CBF-based dynamics, corresponding to an increasing sequence $\{\alpha_{i}\}_{i\in\mathbb{N}}$ of parameters $\alpha_{i}$ with $\lim_{i\to\infty}\alpha_{i}=\infty$, has a subsequence which converges to a solution of the PDS (7). Moreover, the whole sequence converges to a PDS solution (7), when this solution is unique.
@@
Given Props. IV.1, VIII.1 and Cors. IV.1, IV.2, IV.3, the proof of Thm. IV.1 follows steps similar to [10, Section 4]. $\blacksquare$
Assuming that solutions of CBF-based dynamics (10) stay in $\mathcal{Z}\times\mathcal{S}$, Thm. IV.1 establishes their uniform convergence to solutions of PDSs (7). While it is guaranteed that $\xi$ stays in $\mathcal{S}$, this is not always the case for $\zeta$. Nevertheless, if PDS (7) obeys a stability property, then, as the result below shows, solutions to (10) stay in $\mathcal{Z}\times\mathcal{S}$, for sufficiently large $\alpha$, and thus converge to solutions of PDS (7).
@@
Consider vector fields $f:^{m}\times^{n}\to^{n}$ and $g:^{m}\times^{n}\to^{m}$, a compact set $\mathcal{Z}\subset^{m}$ and a set $\mathcal{S}\subset^{n}$ defined by (1). Let Assumptions 1 and 2 hold. Further, let $g$ be continuous and locally bounded. Given $\gamma>0$, consider compact sets $\mathcal{Z}_{0},\mathcal{Z}^{\prime}$, such that $\mathcal{Z}_{0}\subseteq\mathcal{Z}^{\prime}$ and $\mathcal{Z}^{\prime}+\gamma\mathbb{B}\subseteq\mathcal{Z}$. Given any initial condition $(z_{0},x_{0})\in\mathcal{Z}_{0}\times\mathcal{S}$ and a $T>0$, assume any solution $\phi:[0,T]\to^{m}\times^{n}$ to PDS (7), with $\phi(0)=(z_{0},x_{0})$, satisfies $\phi(t)\in\mathcal{Z}^{\prime}\times\mathcal{S}$ for all $t\in[0,T]$. Then, there exists $\alpha^{\prime}\geq\alpha_{*}$, such that for any $\alpha\geq\alpha^{\prime}$, the solution $\psi:[0,T]\to^{m}\times^{n}$ to the CBF-based dynamics (10), with $\psi(0)=(z_{0},x_{0})$, satisfies $\psi(t)\in\mathcal{Z}\times\mathcal{S}$ for all $t\in[0,T]$. Thus, for any increasing sequence $\{\alpha_{i}\}_{i\in\mathbb{N}}$, such that $\alpha_{i}\geq\alpha^{\prime}$ and $\lim_{i\to\infty}\alpha_{i}=\infty$, and respective sequence of solutions $\{\psi_{i}\}_{i\in\mathbb{N}}$ to (10), statements 1 and 2 of Thm. IV.1 hold.
@@
The stability assumption of Thm. IV.2 is often met in practice, as PDS-based controllers [8, 10, 5, 9, 7, 11] normally guarantee closed-loop stability of (7). Thus, if the PDS-based controller is replaced by a CBF-based one, to yield the CBF-based dynamics (10), Thm. IV.2 ensures that trajectories of (10) converge to trajectories of (7); that is, (10) behaves similarly to (7). $\blacksquare$
Remark 6.
We study solutions that are constrained in the compact set $\mathcal{Z}\times\mathcal{S}$. As $\mathcal{Z}\times\mathcal{S}$ can be arbitrarily large, our results are semiglobal. Adopting the state-controller interconnection interpretation of (7) and (10), the set $\mathcal{S}$ describes constraints on controller states, whereas $\mathcal{Z}$ is where state trajectories lie. The larger $\mathcal{Z}$ is, the larger $\alpha_{*}$ is (Prop. IV.1). Thus, for smaller $\mathcal{Z}$, one may use smaller values of $\alpha$, to approximate PDS-based dynamics with CBF-based ones. $\blacksquare$
Remark 7.
As our results are semiglobal, one might be able to show non-uniform convergence in non-compact domains, through limit arguments over sequences of compact sets converging to n. $\blacksquare$
Remark 8.
Combining Theorem IV.2 and [25, Prop. 6.34], one can infer stability of CBF-based dynamics: if the origin for PDS-based dynamics (7) is asymptotically stable, then it is practically stable for CBF-based dynamics (10). Stability and contractivity results for CBF-based dynamics have also been reported in [21]. $\blacksquare$
V Numerical Examples
V-A Feedback optimization
(a) Evolution of the control input
(b) Evolution of state $\zeta_{1}$
(c) Evolution of state $\zeta_{2}$
@@ V-A Feedback optimization
$A=\begin{bmatrix}-1&1\\ 0&-2\end{bmatrix},\text{ }B=\begin{bmatrix}0\\ 1\end{bmatrix},\text{ }\Phi(\zeta)=\Big{(}\zeta-\begin{bmatrix}1\\ 1\end{bmatrix}\Big{)}^{2},\text{ }\mathcal{S}=\{u\in:0.36-u^{2}\geq 0\}$
@@
V-B Control of a synchronverter
(a) Evolution of state $\zeta_{1}=i_{d}$
(b) Evolution of state $\zeta_{2}=i_{q}$
(c) Evolution of state $\zeta_{3}=\omega$
(d) Evolution of state $\zeta_{4}=\delta$
(e) Evolution of output $y_{1}=P$
(f) Evolution of output $y_{2}=Q$
Figure 3: Evolution of the state $\zeta(t)$ and outputs $P,Q$, for controllers with PDS-based dynamics and CBF-based ones, for different $\alpha$.
Figure 4: Trajectories $\xi(t)$ of the dynamic controller with PDS-based dynamics and CBF-based ones, for different $\alpha$.
Synchronverters are inverters that behave like synchronous generators (see [8]). The synchronverter’s state is $\zeta=\begin{bmatrix}i_{d}&i_{q}&\omega&\delta\end{bmatrix}^{\top}$, where $i_{d}$ and $i_{q}$ are the stator currents’ $d$ and $q$ components, $\omega$ is the (virtual) rotor angular velocity, and $\delta$ is the power angle. The control input is $u=\begin{bmatrix}T_{m}&i_{f}\end{bmatrix}^{\top}$, where $T_{m}$ is the (virtual) prime mover torque and $i_{f}$ is the (virtual) field current. The $\zeta$-dynamics is
$\dot{\zeta}=H^{-1}A(\zeta,u)\zeta+H^{-1}v(\zeta,u),\quad y=G(\zeta)$ (16)
where $y=\begin{bmatrix}P&Q\end{bmatrix}^{\top}$ is the active power $P$ and reactive power $Q$,
$H=\begin{bmatrix}L&0&0&0\\ 0&L&0&0\\ 0&0&J&0\\ 0&0&0&1\end{bmatrix},\text{ }A(\zeta,u)=\begin{bmatrix}-R&\omega L&0&0\\ -\omega L&-R&-mi_{f}&0\\ 0&-mi_{f}&-D_{p}&0\\ 0&0&1&0\end{bmatrix},$
$v(\zeta,u)=\begin{bmatrix}V\sin\delta\\ V\cos\delta\\ T_{m}+D_{p}\omega_{n}\\ -\omega_{g}\end{bmatrix},\text{ }G(\zeta)=-V\begin{bmatrix}\cos\delta&\sin\delta\\ -\sin\delta&\cos\delta\end{bmatrix}\begin{bmatrix}i_{q}\\ i_{d}\end{bmatrix}$
[…]
The system admits a steady-state input-output map (see [8]). The control problem is to regulate the output $y$ to a reference $r$. In [8], a PDS-like dynamic controller is proposed, for antiwindup control:
$\dot{\xi}=\Pi_{\mathcal{S}}(\xi,r-y),\quad u=\mathcal{N}(\xi)$ (17)
where the constraint set $\mathcal{S}$ and the map $\mathcal{N}$ are to be designed.
[…]
@@ VI Conclusion
We have proven that trajectories of CBF-based dynamics uniformly converge to those of PDS-based ones, as the parameter $\alpha$ of the CBF constraint approaches $\infty$. We have also proven that CBF-based dynamics are perturbations of PDSs, and quantified the perturbation. One can use our results to implement discontinuous PDS-based controllers in a continuous manner, employing CBFs. We have demonstrated this on a synchronverter control example. Our results also enable novel numerical simulation schemes for PDSs, through CBF-based dynamics, overcoming disadvantages of existing ones, such as projections to possibly non-convex sets. Finally, this bridge between PDSs and CBFs might yield other results in the future, such as insights on stability.
VII Acknowledgements
The authors would like to thank Dr. Pietro Lorenzetti for his invaluable help on the implementation of the synchronverter numerical example.
VIII Technical results and proofs
VIII-A Proof of Prop. IV.1
To prove Prop. IV.1, we use Lemmas VIII.1, VIII.2 and VIII.3 below, which have already appeared in [19], in a simpler form. Here they are extended, to account for interconnections (8) and (10), rather than autonomous setups (6) and (9), which was the context of [19]. [^5] Here, $f:^{m}\times^{n}\to^{n}$, that is $f$ has both $\xi$ and $\zeta$ as arguments. Moreover, in [19] the proofs of Lemmas VIII.1 and VIII.2 were only sketched. Here, we provide the full proofs of Lemmas VIII.1 and VIII.2 and recall Lemma VIII.3. The following results are stated under the assumptions of Prop. IV.1. Given $z\in\mathcal{Z}$, denote $U_{\mathrm{cbf},\alpha}(z):=\{x\in\mathcal{S}\mid\mathcal{L}_{f}h(z,x)+\alpha h(x)\leq 0\}$.
Lemma VIII.1.
Given $\alpha>0$ and $z\in\mathcal{Z}$, consider $x\in U_{\mathrm{cbf},\alpha}(z)$ and $y\in\mathrm{proj}_{\partial\mathcal{S}}(x)$. It holds that $\|x-y\|\leq\gamma(\frac{1}{\alpha}|\mathcal{L}_{f}h(z,x)|)$.
Proof of Lemma VIII.1.